An investment club invested a part of $14,000 in a 9% annual simple interest account and the remainder in a 15% annual simple interest account. The amount of interest earned for one year was $1800. How much was invested in the 9% account?
Answer by richwmiller(17219)

Total amount of money invested: $14000
x+y=14000,
Total yearly interest for the two accounts is: $1800
0.09*x+0.15*y=1800
x=14000-y
Substitute for x
0.09*(14000-y)+0.15*y=1800
Multiply out
1260-0.09*y+0.15*y=1800
Combine like terms.
0.06*y=540
Isolate y
y=$ 9000.00 at 15% earns $1350 interest
x=14000-y
Calculate x
x=$ 5000.00 at 9% earns $450 interest
Check
0.09*5000+0.15*9000=1800
450+1350=1800
1800=1800
If 1800=1800 is TRUE and neither x nor y is negative then all is well
codeint1